OpenStudy (anonymous):

Plz help! ... A system of linear inequalities is shown below: y – x < 0 x – 1 > 0 Which of the following graphs best represents the solution set to this system of linear inequalities?

OpenStudy (anonymous):

graph?

OpenStudy (anonymous):

wait one sec ha

OpenStudy (e.mccormick):

You can test points, or just graph it with a tool... https://www.desmos.com/calculator

OpenStudy (anonymous):

ohhh desmos!!!!

OpenStudy (anonymous):

first one is answers A. and B. second one are answers C. and D.

OpenStudy (anonymous):

ha I just need help with the Q plz

OpenStudy (anonymous):

https://www.desmos.com/calculator/kgqrmf6ewm

Nnesha (nnesha):

but some school teachers don't allow students to use desmos or mathwy.com calculator first step is to solve for y

OpenStudy (anonymous):

what is that for?

Nnesha (nnesha):

\[\huge\rm y - x < 0\] to graph this first you should solve for y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are u a teacher @Nnesha ?

Nnesha (nnesha):

nope. solve that for y

OpenStudy (anonymous):

ha I see that but thats just a name

OpenStudy (anonymous):

and ok

Nnesha (nnesha):

do you know how to solve for y ??\[y - x<0\]

OpenStudy (anonymous):

y= 0 + x?

OpenStudy (anonymous):

idk

Nnesha (nnesha):

yes right but there is inequality sign

OpenStudy (anonymous):

and u don't need that 0. or that + sign.

Nnesha (nnesha):

\[\huge\rm y < x\] x + 0 =x so

OpenStudy (anonymous):

oh ok so what next?

Nnesha (nnesha):

solve 2nd inequity for x

OpenStudy (anonymous):

so x= 0 - x?

OpenStudy (anonymous):

idk im not good at math

Nnesha (nnesha):

nope \[\huge\rm x - 1 >0\] you have to move -1 to the left side

OpenStudy (anonymous):

oh so x= 0 + 1?

Nnesha (nnesha):

yep right but again ">" inequality sign

OpenStudy (anonymous):

oh yes sorry

OpenStudy (anonymous):

so do u know the answer or what else do we need to do?

Nnesha (nnesha):

\[\huge\rm y < x \] \[\huge\rm x > 1 \] now you can graph it ">'" "<" less than or greater sign so you should draw dotted line y = mx + b slope intercept form equation where m is slope and b is y-intercept x = something vertical line

OpenStudy (anonymous):

sorry but i still dont know how to get the answer

OpenStudy (anonymous):

i don't need to graph any thing my the way, it;s multiple chose.

Nnesha (nnesha):

shade above the line \[\huge\rm > \] \[\huge\rm < \] shade below the line

OpenStudy (anonymous):

u need to graph it in order to know which graph u need to choose

Nnesha (nnesha):

look at ur answer choices which graph have vertical line ??

OpenStudy (anonymous):

ohh ok

Nnesha (nnesha):

\[\huge\rm x > 1\] vertical line shade right side

OpenStudy (anonymous):

they all have vertical lines so is the answer D.?

OpenStudy (e.mccormick):

And as I said, you can do test points. When you have 4 answers given, sometimes just picking a few points, plugging them in, and seeing if they are true is a good way to eliminate things.

OpenStudy (anonymous):

yeah that's true @e.mccormick

Nnesha (nnesha):

 x > 1 vertical line shade right side